A 0.5 kg ball is thrown upwards with an initial velocity of 6 m/s. How high will the ball reach before it starts falling again?
1
Expert's answer
2020-05-18T10:10:29-0400

Equations of motion of the ball are "v(t) = v_0 - g t" and "y(t) = v_0 t - \frac{g t^2}{2}".

When the ball reaches its highest point, the speed is zero, hence "v(t_s) = 0 \Rightarrow t_s = \frac{v_0}{g}".

"t_s" is the time it took to reach the highest point. Therefore, "y" coordinate at that moment of time is how high will the ball reach before it starts falling again: "H = y(t_s) = \frac{v_0^2}{g} - \frac{v_0^2}{2 g} = \frac{v_0^2}{2 g} \approx 1.83 m".